OpenStudy (anonymous):

A platform has a volume of 135 cubic feet, but the concrete Joey wants to buy is sold by the cubic yard. How many cubic yards of concrete does Joey need to buy?

OpenStudy (anonymous):

a yard is equal to 3 feet so you need to divide by 3 to the power of 3 because its cubic, so you need: 135 / 3^3 = 135 / 27 = 5 Joey needs to buy 5 cubic yards of concrete.

OpenStudy (anonymous):

Because a yard is equal to 3 feet then you can imagine a cube made up of little cubes that are all 1 foot high/wide/ and deep. This cube will be three cubes high, three cubes wide and three cubes deep. It will have 27 or \[3^{3}\] little cubes in it. Now you know 1 cubic yard = 27 cubic feet. You have a ratio \[1 (yd)^{3}/27 (ft)^{3}\]. Then you can use this to convert \[135 (ft)^{3}\times1 (yd)^{3}/27 (ft)^{3}=5 (yd)^{3}\]